I need to know the answer to 1 1/15 divided by 1 19/20 thank you in advance for your help
Answer by Alan3354(23858)

I need to know the answer to 1 1/15 divided by 1 19/20 thank you in advance for your help
-----------
Step 1 - change to fractions
1 1/15 = 16/15
1 19/20 = 39/20
-----
Step 2 - invert the divisor: 39/20 --> 20/39
----
Step 3, multiply
= (16/15)*(20/39)
Step 4 - do it